OpenStudy (anonymous):

What is the equation of the line that passes through (-3, -1) and has a slope of 3/5? Put your answer in slope-intercept form.

OpenStudy (anonymous):

(y-y1)=m(x-x1) (y-(-1))=3/5(x-(-3)) y+1=3/5x+9/5 y=3/5x+9/5(-1) y=3/5x+4/5